# Conceptualization of this study, Methodology, Software

### v1

@@ 4.3.4 Denoising effect of each method at all noise levels
Table 3 illustrates the denoising efficacy of various methods on EOG and EMG datasets. The outcomes in this table lead to the following: insights:
@@
    Featuring a more intricate architecture incorporating diverse convolutional kernels for multi-scale feature extraction and alleviating overfitting through the introduction of residual connections, 1D-ResCNN surpasses SCNN, significantly enhancing denoising outcomes.
@@
    Leveraging the transformer architecture, EEGDnet excels in denoising, benefitting from the global modeling prowess of the attention mechanism, complemented by residual connections and layer normalization. This results in substantial denoising improvements compared to SCNN and 1D-ResCNN.
@@
    Capitalizing on the Retnet framework, EEGDiR achieves superior denoising performance by comprehensively understanding input temporal information and exhibiting robust global modeling capabilities. The incorporation of residuals and multiple normalizations (layer norm, group norm) further distinguishes EEGDiR, outperforming other networks. Moreover, guided by our proposed signal embedding, EEGDiR intelligently processes temporal information, forming patches from consecutive samples. This strategy adeptly captures the contextual and temporal relationships within EEG signals, aligning with their prolonged temporal characteristics. The judicious embedding strategy contributes to optimized denoising performance, reinforcing EEGDiR’s exceptional superiority over alternative networks.
Table 3: Average performances of all SNRs (from $-7$ dB to $2$ dB). The smaller $RRMSE_{temporal}$ and $RRMSE_{spectral}$ , and the larger $CC$, the better denoising effect. Note that all the models are trained and tested on the same data set. The baseline of EEGDiR consists of $4$ layers and $8$ Heads with patch size $16$ and hidden dim $512$. For $RRMSE_{temporal}$ , $RRMSE_{spectral}$ , the lower the better. For $CC$, the higher the better. The best result is shown in bold.
Model Ocular artifact Muscle artifact
$RRMSE_{temporal}$ $RRMSE_{spectral}$ $CC$ $RRMSE_{temporal}$ $RRMSE_{spectral}$ $CC$
SCNN 0.6176 0.5905 0.7938 0.7342 0.7977 0.7364
1D-ResCNN 0.5409 0.5900 0.8503 0.6921 0.6848 0.7434
EEFDnet 0.4819 0.4647 0.8725 0.6200 0.5565 0.7711
EEGDiR(ours) 0.3279 0.3616 0.9329 0.5322 0.5004 0.8072
@@ 4.3.5 Denoising effect of each method at different noise levels
In the subsequent section, we present the quantitative benchmarking results ($RRMSE_{temporal}$, $RRMSE_{spectral}$, $CC$ ) of diverse methods across varying SNR levels in the test set. Figures 3 and 4 showcase the test outcomes on the EOG and EMG test sets, pivotal for evaluating the denoising efficacy of the methods.
@@
    Among the methods, SCNN displays the highest $RRMSE_{temporal}$ and $RRMSE_{spectral}$, along with the lowest $CC$. This indicates SCNN’s inferior denoising performance, attributed to its relatively simple network structure hindering effective input feature extraction. In contrast, the more intricate 1D-ResCNN yields significantly improved denoising outcomes. However, compared to EEGDnet with a Transformer model and global modeling capability, there are discernible performance gaps. The EEGDiR model, incorporating Retnet, achieves the lowest $RRMSE_{temporal}$ and $RRMSE_{spectral}$, coupled with the highest $CC$. It excels in denoising tasks across varying noise levels.
@@
    Analyzing the $RRMSE_{temporal}$ results on the EOG dataset, denoising performance improves with decreasing noise levels and increasing SNR levels across all methods. However, the performance gap between methods persists, potentially due to EOG noise being more easily removed than EMG noise. On the EMG dataset, the performance gap diminishes as noise levels decrease (SNR levels increase), particularly evident for SCNN, 1D-ResCNN, and EEGDnet. Nevertheless, EEGDiR maintains superior denoising performance.
@@
    Evaluation of the $RRMSE_{spectral}$ results on the EOG dataset indicates weaker denoising performance for SCNN and 1D-ResCNN, possibly due to limited global modeling capability. Conversely, EEGDnet and EEGDiR exhibit superior denoising performance owing to their robust global modeling ability. On the EMG dataset, despite decreasing differences in performance as noise levels decrease, EEGDnet and EEGDiR consistently outperform SCNN and 1D-ResCNN.
@@
    Examination of $CC$ results on the EOG dataset reveals improved denoising performance for all methods as noise levels decrease, with relatively stable performance differences. In the EMG dataset, SCNN exhibits poorer performance due to the dataset’s more complex noise. Conversely, the denoising performance of the remaining three networks improves as noise levels decrease, with consistent performance differences. Notably, EEGDiR maintains excellent denoising performance throughout.
@@
Figure 3: Performance of four deep-learning networks at different SNR levels with ocular artifact removal. The smaller $RRMSE_{temporal}$ and $RRMSE_{spectral}$ , and the larger Correlation Coefficient($CC$), the better denoising effect. The denoising performance increases as the SNR increases.
@@
Figure 4: Performance of four deep-learning networks at different SNR levels with muscle artifact removal. The smaller $RRMSE_{temporal}$ and $RRMSE_{spectral}$, and the larger Correlation Coefficient($CC$), the better denoising effect. The denoising performance increases as the SNR increases.
@@
Figure 5: Performance of four DL networks (SCNN, 1D-ResDNN, EEGDnet, EEGDiR) in ocular artifact removal. The smaller $RRMSE_{temporal}$ and $RRMSE_{spectral}$, and the larger Correlation Coefficient($CC$), the better denoising effect. EEGDiR models robustly outperform other model for EEG denosing.
@@
Figure 6: Performance of four DL networks (SCNN, 1D-ResDNN, EEGDnet, EEGDiR) in muscle artifact removal. The smaller $RRMSE_{temporal}$ and $RRMSE_{spectral}$, and the larger Correlation Coefficient($CC$), the better denoising effect. EEGDiR models robustly outperform other model for EEG denosing.
(a) Denoising outcomes on Ocular Artifact.
(b) Denoising outcomes on Muscle Artifact.
Figure 7: Visualization of denoising outcomes for various state-of-the-art models: (a) Denoising outcomes on Ocular Artifact. (b) Denoising outcomes on Muscle Artifact. A closer inspection can be facilitated by zooming in for a more detailed view. It is important to highlight that we have restored the network output’s amplitude back to the original data scale through back-normalization. The temporal domain operates at a sampling rate of 256 SPS. From the provided results, it is evident that the denoising results achieved by the proposed EEGDiR model in this study closely approximate the true signal.
Figures 5 and 6 illustrate the ANOVA results for models evaluated on the EOG and EMG datasets. Drawing conclusions from the provided information and ANOVA analyses, the following observations emerge:
  1. [(1)]
  2. 1.
    $RRMSE_{temporal}$ Rancking: The denoising performance across the four methods is observed as follows: CNN < 1D-ResCNN < EEGDnet < EEGDiR. ANOVA analysis indicates significant differences in $RRMSE_{temporal}$, with marked distinctions in the EOG dataset and relatively modest differences in the EMG dataset. EEGDiR significantly outperforms other methods in time-domain denoising for both EOG and EMG datasets, followed by EEGDnet and 1D-ResCNN, while SCNN exhibits the least efficacy.
  3. 2.
    $RRMSE_{spectral}$ Rancking: The denoising performance order for $RRMSE_{spectral}$ is 1D-ResCNN < SCNN < EEGDnet < EEGDiR. The occurrence of 1D-ResCNN < SCNN is attributed to 1D-ResCNN’s superior extraction of features in the time domain, leading to diminished denoising performance in the spectral features. ANOVA results show a significant difference in $RRMSE_{spectral}$ among methods on the EOG dataset, while the difference is relatively weak on the EMG dataset. EEGDiR significantly outperforms other methods in spectral denoising, followed by EEGDnet and SCNN, while 1D-ResCNN is less effective.
  4. 3.
    $CC$ Metric Ranking: The denoising performance sequence on the $CC$ metric remains SCNN < 1D-ResCNN < EEGDnet < EEGDiR. ANOVA analysis reveals a significant difference in $CC$ between methods for the EOG dataset, while the difference is relatively weak for the EMG dataset. Comparing mean values, EEGDiR excels in correlation, followed by EEGDnet and 1D-ResCNN, while SCNN exhibits poor CC performance.
In summary, the outstanding denoising performance of the EEGDiR method can be attributed to multiple factors. The Retnet architecture provides enhanced global modeling capability, enabling a more accurate restoration of input timing information. The proposed signal embedding method adeptly handles the prolonged temporal information of EEG signals, capturing context and temporal relationships intelligently through the combination of successive sampling points into patches. This advantage enables EEGDiR to achieve superior denoising effects in both the time domain and spectral characteristics. Additionally, the synergy of residual connectivity and multiple normalization methods (layer norm, group norm) enhances EEGDiR’s denoising performance and robustness to noise. The advanced Retnet architecture, skillful embedding strategy, and enhanced network design collectively contribute to EEGDiR’s exceptional performance in time-domain and spectral denoising, as well as correlation.
The visualization results depicting the impact of EOG and EMG noise on EEG signals are presented in Figure 7a and Figure 7b, yielding the following observations. It is noteworthy that the dataset has undergone variance normalization. When presenting the visualization results, Equation (15) is employed to scale down the results to the original data scale, enhancing the accuracy of showcasing the noise effect on EEG signals.
@@ 4.3.6 Visualization of denoising results for each method on EOG and EMG datasets
    The relatively complex structure and residual connectivity of 1D-ResCNN result in an improvement in denoising compared to SCNN, emphasizing the impact of network architecture on denoising performance. Leveraging the global modeling and feature extraction capabilities facilitated by the Transformer’s attention mechanism, EEGDnet outperforms SCNN and 1D-ResCNN in denoising. This underscores the enhancement of network performance with the introduction of the attention mechanism.
@@
    Overall, the denoising effect achieved by EEGDiR closely approaches that of a noise-free signal. This notable advantage can be attributed to the synergistic effect of the Retnet architecture and our proposed signal embedding, tailored to match the characteristics of EEG signals. Firstly, the Retnet architecture enhances understanding of timing information in EEG signals through its robust global modeling capability. This enables the network to accurately capture the complex time-domain structure, thereby improving denoising performance. Secondly, our signal embedding method adeptly addresses the challenge of handling long temporal information in EEG signals. By intelligently grouping consecutive sampling points into patches, this method effectively preserves the context and temporal relationships of the signal, facilitating the network in learning and restoring features more efficiently. The sensitivity to long temporal information aligns with the characteristics of EEG signals, forming the basis for EEGDiR’s outstanding denoising effect. Therefore, the performance of EEGDiR in approaching noise-free signals arises not only from the superior processing of temporal information by the Retention mechanism but also from the mutually reinforcing capabilities of Retnet and signal embedding. This synergy enables the network to better comprehend and process the intricate structure of EEG signals.
@@ 5 Discussions
In contrast to prevailing deep learning denoising methodologies, our approach involves incorporating the Retnet architecture into Long Short-Term Memory (LLM) for EEG signal denoising. The Retnet architecture exhibits notable strengths in comprehending temporal information and excelling in global modeling feature extraction. The decision to integrate Retnet stems from its proficiency in handling temporal information and its exceptional global modeling performance. Concurrently, we introduce signal embedding to intelligently manage the protracted time-series information inherent in EEG signals. This approach enhances context and temporal relationship preservation by aggregating consecutive sampling points into patches. Given the prolonged time-series nature of EEG signals, this method aligns well with their characteristics, contributing to enhanced denoising performance. Hence, the incorporation of Retnet and patch embedding serves not only to refine the comprehension and processing of EEG signals’ temporal structure but also to augment denoising efficacy from diverse perspectives. This dual strategy of leveraging Retnet and patch embedding renders our method adaptable to the intricate characteristics of EEG signals, resulting in superior denoising outcomes.
Retnet’s adept handling of temporal information in processing EEG signals extends its applicability to various temporal signals, including electromagnetic, seismic, and biomedical signals. Its affinity for temporal information bestows a distinct advantage when addressing signals necessitating comprehensive time-based modeling. This advantageous trait is particularly pertinent to other temporal signals like electromagnetic, seismic, and biomedical signals. The friendly temporal information processing of Retnet proves beneficial in scenarios requiring a global temporal perspective. The incorporation of signal embedding in EEG signals enhances the preservation of contextual and temporal relationships within the signals. This characteristic, valuable for signal processing across various domains, especially in cases with extensive temporal information or a requirement for comprehensive temporal modeling, extends beyond EEG signals. The network proposed in this paper achieves notable performance enhancements by amalgamating Retnet’s global modeling and signal embedding’s temporal information processing strategies. The broad applicability of this strategy suggests its potential effectiveness in diverse signal processing domains, particularly tasks demanding a synthesis of global and temporal information. The integrated strategy of Retnet and signal embedding may evolve into a generalized signal processing approach, particularly suitable for scenarios requiring consideration of both global features and temporal information. Subsequent research endeavors could explore the application of this approach to other signal types, validating its adaptability and performance across distinct domains.

### v2

@@ 4.3.4 Denoising effect of each method at all noise levels
Table 3 illustrates the denoising efficacy of various methods on EOG, EMG and SS2016 EOG datasets. The outcomes in this table lead to the following:
@@
    Featuring a more intricate architecture incorporating diverse convolutional kernels for multi-scale feature extraction and alleviating overfitting through the introduction of residual connections, 1D-ResCNN surpasses SCNN, significantly enhancing denoising outcomes. Thanks to the temporal information added to the input by the LSTM structure, LSTM exhibits better denoising results than 1D-ResCNN.
@@
    Leveraging the transformer architecture, EEGDnet excels in denoising, benefitting from the global modeling prowess of the attention mechanism, complemented by residual connections and layer normalization. This results in substantial denoising improvements compared to SCNN, 1D-ResCNN and LSTM.
@@
    Capitalizing on the Retentive Network, EEGDiR achieves superior denoising performance by comprehensively understanding input temporal information and exhibiting robust global modeling capabilities. The incorporation of residuals and multiple normalizations (layer norm, group norm) further distinguishes EEGDiR, outperforming other networks. Moreover, guided by our proposed signal embedding, EEGDiR intelligently processes temporal information. This strategy adeptly captures the contextual and temporal relationships within EEG signals, aligning with their prolonged temporal characteristics. The signal embedding strategy contributes to optimized denoising performance, reinforcing EEGDiR’s exceptional superiority over alternative networks.
Table 3: Average performances of all SNRs (from $-7$ dB to $2$ dB). The smaller $RRMSE_{t}$ and $RRMSE_{s}$ , and the larger $CC$, the better denoising effect. Note that all the models are trained and tested on the same data set. The baseline of EEGDiR consists of $4$ layers and $8$ Heads with patch size $16$ and hidden dim $512$. For $RRMSE_{t}$ , $RRMSE_{s}$ , the lower the better. For $CC$, the higher the better. The best result is shown in bold.
Model EOG dataset EMG dataset SS2016 EOG dataset
$RRMSE_{t}$ $RRMSE_{s}$ $CC$ $RRMSE_{t}$ $RRMSE_{s}$ $CC$ $RRMSE_{t}$ $RRMSE_{s}$ $CC$
SCNN 0.6176 0.5905 0.7938 0.7342 0.7977 0.7364 0.5893 0.6724 0.8156
1D-ResCNN 0.5409 0.5900 0.8503 0.6921 0.6848 0.7434 0.5523 0.5804 0.8552
LSTM 0.5290 0.4894 0.8449 0.6560 0.6092 0.7461 0.4823 0.5573 0.8747
EEGDnet 0.4819 0.4647 0.8725 0.6200 0.5565 0.7711 0.4594 0.5267 0.8875
EEGDiR(ours) 0.3279 0.3616 0.9329 0.5322 0.5004 0.8072 0.3146 0.3613 0.9488
@@ 4.3.5 Denoising effect of each method at different noise levels
In the subsequent section, we present the quantitative benchmarking results ($RRMSE_{t}$, $RRMSE_{s}$, $CC$ ) of diverse methods across varying SNR levels in the test set. Figures 5, Figures 6 and Figures 7 showcase the test outcomes on the EOG, EMG and SS2016 EOG test dataset, pivotal for evaluating the denoising efficacy of the methods.
@@
    Among the methods, SCNN displays the highest $RRMSE_{t}$ and $RRMSE_{s}$, along with the lowest $CC$. This indicates SCNN inferior denoising performance, attributed to its relatively simple network structure hindering effective input feature extraction. In contrast, the more intricate 1D-ResCNN and LSTM yields significantly improved denoising outcomes. However, compared to EEGDnet with a Transformer model and global modeling capability, there are discernible performance gaps. The EEGDiR, incorporating Rententive Network and signal embedding, achieves the lowest $RRMSE_{t}$ and $RRMSE_{s}$, coupled with the highest $CC$. It excels in denoising tasks across varying noise levels.
@@
    Analyzing the $RRMSE_{t}$ results on the EOG and SS2016 EOG dataset, denoising performance improves with decreasing noise levels and increasing SNR levels across all methods. However, the performance gap between methods persists, potentially due to EOG noise being more easily removed than EMG noise. On the EMG dataset, the performance gap diminishes as noise levels decrease (SNR levels increase), particularly evident for SCNN, 1D-ResCNN, LSTM, and EEGDnet. Nevertheless, EEGDiR maintains superior denoising performance.
@@
    Evaluation of the $RRMSE_{s}$ results on the EOG and SS2016 EOG dataset indicates weaker denoising performance for SCNN, 1D-ResCNN and LSTM, possibly due to limited global modeling capability. Conversely, EEGDnet and EEGDiR exhibit superior denoising performance owing to their robust global modeling ability. On the EMG dataset, despite decreasing differences in performance as noise levels decrease, EEGDnet and EEGDiR consistently outperform SCNN, 1D-ResCNN and LSTM.
@@
    Examination of $CC$ results on the EOG and SS2016 EOG dataset reveals improved denoising performance for all methods as noise levels decrease, with relatively stable performance differences. In the EMG dataset, SCNN exhibits poorer performance due to the dataset’s more complex noise. Conversely, the denoising performance of the remaining three networks improves as noise levels decrease, with consistent performance differences. Notably, EEGDiR maintains excellent denoising performance throughout.
@@
Figure 5: Performance of four deep-learning networks at different SNR levels with EOG dataset artifact removal. The smaller $RRMSE_{t}$ and $RRMSE_{s}$ , and the larger Correlation Coefficient($CC$), the better denoising effect. The denoising performance increases as the SNR increases.
@@
Figure 6: Performance of four deep-learning networks at different SNR levels with EMG dataset artifact removal. The smaller $RRMSE_{t}$ and $RRMSE_{s}$, and the larger Correlation Coefficient($CC$), the better denoising effect. The denoising performance increases as the SNR increases.
@@
Figure 7: Performance of four deep-learning networks at different SNR levels with SS2016 EOG dataset artifact removal. The smaller $RRMSE_{t}$ and $RRMSE_{s}$, and the larger Correlation Coefficient($CC$), the better denoising effect. The denoising performance increases as the SNR increases.
Figures 8, Figures 9 and Figures 10 illustrate the ANOVA results for models evaluated on the EOG, EMG and SS2016 EOG datasets. Drawing conclusions from the provided information and ANOVA analyses, the following observations emerge:
  1. [(1)]
  2. 1.
    $RRMSE_{t}$ Rancking: The denoising performance across the four methods is observed as follows: SCNN < 1D-ResCNN < LSTM < EEGDnet < EEGDiR. ANOVA analysis indicates significant differences in $RRMSE_{t}$, with marked distinctions in the EOG dataset and relatively modest differences in the EMG dataset. EEGDiR significantly outperforms other methods in time-domain denoising for both EOG, EMG and SS2016 EOG datasets, followed by EEGDnet, LSTM and 1D-ResCNN, while SCNN exhibits the least efficacy.
  3. 2.
    $RRMSE_{s}$ Rancking: The denoising performance order for $RRMSE_{s}$ is 1D-ResCNN < SCNN < LSTM < EEGDnet < EEGDiR. The occurrence of 1D-ResCNN < SCNN is attributed to 1D-ResCNN superior extraction of features in the time domain, leading to diminished denoising performance in the spectral features. ANOVA results show a significant difference in $RRMSE_{s}$ among methods on the EOG and SS2016 EOG dataset, while the difference is relatively weak on the EMG dataset. EEGDiR significantly outperforms other methods in spectral denoising, followed by EEGDne, LSTM and SCNN, while 1D-ResCNN is less effective.
  4. 3.
    $CC$ Metric Ranking: The denoising performance sequence on the $CC$ metric remains SCNN < 1D-ResCNN < LSTM < EEGDnet < EEGDiR. ANOVA analysis reveals a significant difference in $CC$ between methods for the EOG and SS2016 EOG dataset, while the difference is relatively weak for the EMG dataset. Comparing mean values, EEGDiR excels in correlation, followed by EEGDnet and 1D-ResCNN, while SCNN exhibits poor CC performance.
In summary, the outstanding denoising performance of the EEGDiR method can be attributed to multiple factors. The Rentetive Network architecture provides enhanced global modeling capability, enabling a more accurate restoration of input timing information. The proposed signal embedding method adeptly handles the prolonged temporal information of EEG signals, capturing context and temporal relationships intelligently through the combination of successive sampling points into patches. This advantage enables EEGDiR to achieve superior denoising effects in both the time domain and spectral characteristics. Additionally, the synergy of residual connectivity and multiple normalization methods (layer norm, group norm) enhances EEGDiR denoising performance and robustness to noise. The advanced Retnet architecture, skillful embedding strategy, and enhanced network design collectively contribute to EEGDiR exceptional performance in time-domain and spectral denoising, as well as correlation.
The visualization results depicting the impact of EOG and EMG noise on EEG signals are presented in Figure 11a, Figure 11b and Figure 11c, yielding the following observations. It is noteworthy that the dataset has undergone variance normalization. When presenting the visualization results, Equation (15) is employed to scale down the results to the original data scale, enhancing the accuracy of showcasing the noise effect on EEG signals.
@@ 4.3.6 Visualization of denoising results for each method on EOG and EMG datasets
Figure 8: Performance of four DL networks (SCNN, 1D-ResDNN, LSTM, EEGDnet, EEGDiR) in EOG dataset artifact removal. The smaller $RRMSE_{t}$ and $RRMSE_{s}$, and the larger Correlation Coefficient($CC$), the better denoising effect. EEGDiR models robustly outperform other model for EEG denosing.
(a) $RRMSE_{temporal}$
(b) $RRMSE_{spectral}$
(c) $CC$
Figure 9: Performance of four DL networks (SCNN, 1D-ResDNN, LSTM, EEGDnet, EEGDiR) in EMG dataset artifact removal. The smaller $RRMSE_{t}$ and $RRMSE_{s}$, and the larger Correlation Coefficient($CC$), the better denoising effect. EEGDiR models robustly outperform other model for EEG denosing.
(a) $RRMSE_{temporal}$
(b) $RRMSE_{spectral}$
(c) $CC$
Figure 10: Performance of four DL networks (SCNN, 1D-ResDNN, LSTM, EEGDnet, EEGDiR) in SS2016 EOG dataset artifact removal. The smaller $RRMSE_{temporal}$ and $RRMSE_{spectral}$, and the larger Correlation Coefficient($CC$), the better denoising effect. EEGDiR models robustly outperform other model for EEG denosing.
@@
    The relatively complex structure and residual connectivity of LSTM and 1D-ResCNN result in an improvement in denoising compared to SCNN, emphasizing the impact of network architecture on denoising performance. Leveraging the global modeling and feature extraction capabilities facilitated by the Transformer’s attention mechanism, EEGDnet outperforms SCNN, LSTM and 1D-ResCNN in denoising. This underscores the enhancement of network performance with the introduction of the attention mechanism.
@@
    Overall, the denoising effect achieved by EEGDiR closely approaches that of a noise-free signal. This notable advantage can be attributed to the synergistic effect of the Retentive Network architecture and our proposed signal embedding, tailored to match the characteristics of EEG signals. Firstly, the Retentive Network architecture enhances understanding of timing information in EEG signals through its robust global modeling capability. This enables the network to accurately capture the complex time-domain structure, thereby improving denoising performance. Secondly, our signal embedding method adeptly addresses the challenge of handling long temporal information in EEG signals. By intelligently grouping consecutive sampling points into patches, this method effectively preserves the context and temporal relationships of the signal, facilitating the network in learning and restoring features more efficiently. The sensitivity to long temporal information aligns with the characteristics of EEG signals, forming the basis for EEGDiR outstanding denoising effect. Therefore, the performance of EEGDiR in approaching noise-free signals arises not only from the superior processing of temporal information by the Retention mechanism but also from the mutually reinforcing capabilities of Retentive Network and signal embedding. This synergy enables the network to better comprehend and process the intricate structure of EEG signals.
(a) Denoising outcomes on EOG dataset.
(b) Denoising outcomes on EMG dataset.
(c) Denoising outcomes on SS2016 EOG dataset
Figure 11: Visualization of denoising outcomes for various state-of-the-art models: (a) Denoising outcomes on EOG dataset. (b) Denoising outcomes on EMG dataset. (b) Denoising outcomes on SS2016 EOG dataset. A closer inspection can be facilitated by zooming in for a more detailed view. It is important to highlight that we have restored the network output’s amplitude back to the original data scale through back-normalization. The temporal domain operates at a sampling rate of 256 SPS for EOG and EMG dataset. The temporal domain operates at a sampling rate of 200 SPS for SS2016 EOG dataset. From the provided results, it is evident that the denoising results achieved by the proposed EEGDiR model in this study closely approximate the true signal.
@@ 5 Discussions
In our analysis, we compared various denoising methodologies by examining their efficacy on the EOG, EMG, and SS2016 EOG datasets as presented in Table 3. Our findings underscore the significant advancements our EEGDiR model, incorporating the Retentive Network and signal embedding techniques, makes over other methods.
Comparative Analysis of Denoising Methods. The data reveals that simpler models like SCNN, despite their utility, fall short in more complex noise environments due to their basic structures and lack of advanced features like residual connections. Conversely, 1D-ResCNN improves upon this by utilizing diverse convolutional kernels and residual connections to mitigate overfitting, thereby enhancing denoising results. Notably, the LSTM model, which integrates temporal information directly into its structure, outperforms 1D-ResCNN by better handling the dynamics within EEG signals. EEGDnet, leveraging the global modeling capabilities of the transformer architecture, enhanced by residual connections and layer normalization, significantly surpasses both SCNN and 1D-ResCNN. The most robust performance, however, is exhibited by EEGDiR. This model not only understands complex temporal sequences better due to the Retentive Network but also optimizes the handling of these sequences through our innovative signal embedding approach. This dual strategy is particularly effective in preserving the contextual and temporal integrity of EEG signals, which is crucial given their prolonged time-series nature.
Performance Across Different Noise Levels. Our study also evaluated the performance of these methods across varying SNR levels, as depicted in Figures 5, 6, and 7. All methods demonstrated declining performance with decreasing SNR, highlighting the challenges posed by increased noise levels. However, EEGDiR consistently outperformed other methods at all noise levels, achieving the lowest RRMSE and highest correlation coefficients. This suggests that EEGDiR’s architecture and signal processing strategies are well-suited to effectively reduce noise while preserving the integrity of the EEG signal.
Broader Implications. The success of the Retentive Network in this context not only paves the way for its use in EEG signal denoising but also suggests its applicability to other types of temporal signals, such as electromagnetic and seismic data. The ability of the Retentive Network to process temporal information effectively, coupled with our signal embedding technique, offers a robust framework for denoising tasks across various domains requiring detailed temporal analysis. Moreover, this framework holds significant potential for enhancing downstream EEG tasks, such as classification [19], motor imagery [7], brain recognition [18] and fatigue detection [6], suggesting broad applicability in enhancing the accuracy and effectiveness of these complex applications.